How many moles of N2 reacted if 0.65 mol of NH3 is produced? Please answer in 2 significant figures


1
Expert's answer
2021-06-28T08:15:44-0400

N2 + 3H2 → 2NH3

According to the reaction equation:

"n(N_2) = \frac{1}{2}n(NH_3) = \frac{1}{2} \times 0.65 = 0.32 \;mol"

Answer: 0.32 mol